A small charged bead has a mass of 1.1 g . It is held in a uniform electric field E⃗ = (200,000 N/C, up). When the bead is released, it accelerates upward with an acceleration of 12 m/s2 .What is the charge on the bead? Express your answer with the appropriate units.
Answer #1

Solution)

Since the charge is accelerating upward hence its positive.

Now,

electric force on the bead = q*E = 200000*q N

Also,

Fnet = m*a

200000*q - 1.1*9.81*10^-3 = 1.1*10^-3 * 12

200000*q = 0.02398

q = 1.19*10^-7 C (Ans)

=======

Good luck!:)